# Approximating Feynman Integrals Using Complete Monotonicity and Stieltjes Properties

### v1

@@ 1 Introduction
Feynman integrals are crucial components in the computation of observables in perturbative quantum field theory, in particular for scattering amplitudes relevant in collider physics and gravitational wave physics Travaglini:2022uwo; Huss:2025nlt; Buonanno:2022pgc. In recent years, the development of new mathematical structures has significantly advanced our ability to compute and understand Feynman integrals, see for instance ref. Abreu:2022mfk. However, the full analytic evaluation of multi-loop Feynman integrals remains a challenge, especially when multiple scales are involved. At the same time, there is considerable interest in numerically evaluating observables. It is therefore interesting to ask: How can we leverage analytic insights for numerical evaluation of Feynman integrals?
In the literature, a number of promising methods exist, as we review presently. These include methods based on contour deformation and sector decomposition Binoth:2000ps of Feynman integrals in Feynman parametrization. Sector decomposition helps to disentangle overlapping singularities by reorganizing the integral into a form in which divergences become more transparent and easier to handle numerically. See for refs. Heinrich:2023til; Smirnov:2015mct for computer implementations of this method. The authors of ref. Borinsky:2023jdv use tropical Monte-Carlo-based sampling methods. This method, while powerful and performant, is currently available for finite integrals only.
Another approach is to derive and to numerically evaluate differential equations satisfied by Feynman integrals. For example, refs. Hidding:2020ytt; Armadillo:2022ugh; Prisco:2025wqs; Mezzarobba2025Dynaverse rely on knowing the value of the integrals at some reference point (e.g. as determined by physical consistency conditions, or via other means, such as numerical evaluation via the above methods), and then analytically continue that value to other kinematic values by matching the domains of convergence of series expansions about different kinematic points. For this to work well, it is desirable to have the differential equation in a form as simple as possible (cf. the review Henn:2014qga), which requires additional steps. On the other hand, ref. Liu:2022chg introduces an auxiliary parameter, with the help of which it becomes easy to determine boundary values of an integral, which are then analytically continued to the integral without the auxiliary parameter. While very powerful and universally applicable, introducing the additional scale of the auxiliary parameter sometimes constitutes a bottleneck.
@@
The first approach we consider also takes differential equations as input, but does not require any specific form of them, which makes our approach easy to use and without complicated prerequisites. More specifically, we propose a new method for bootstrapping Feynman integrals, by combining their differential equations with the constraint of complete monotonicity. Complete monotonicity is a powerful positivity condition on (all) derivatives of a function. In ref. Henn:2024qwe two of the present authors showed that many relevant quantities in quantum field theory are completely monotone. See refs. Ma:2025pvq; Mazzucchelli:2025gyg for further recent research in this direction.
In particular, scalar Feynman integrals exhibit complete monotonicity in their Euclidean region with respect to kinematic invariants Henn:2024qwe. When combined with differential equations for a set of integrals, this property imposes powerful constraints on the space of possible solutions. We show that this interplay between differential equations and complete monotonicity can uniquely determine the integral. This extends previous work on numerically bootstrapping Feynman integrals Zeng:2023jek, which employed positivity properties of functions, but did not consider derivatives.
@@
Stieltjes functions enjoy remarkable analyticity properties that imply that they can be extremely well described by Padé approximants Basdevant1972Pade; BGM. In particular, strong bounds and convergence theorems are known for the Padé approximants. Leveraging this knowledge, we show how one can obtain reliable numerical approximations over a wide range of kinematic space, including complex-valued, analytically continued results in physical regions. This allows for a fast and high-precision numerical determination of integrals across a wide range of kinematics by exploiting minimal analytic assumptions to obtain tight numerical constraints.
@@ 2.1 Complete monotonicity and Feynman integrals
A real-valued function $f:R\to\mathbb{R}$ is said to be completely monotonic (CM) in the region $R$ if it satisfies (cf. ref. widder1941laplace)
@@
There is a straightforward connection to Feynman integrals that was uncovered in ref. Henn:2024qwe. In that reference, two of the present authors showed that scalar Feynman integrals are CM functions of appropriate linear combinations of kinematic variables, within their Euclidean region. In this context, we understand by Euclidean region the set of kinematic values for which the graph polynomials appearing in Feynman parametrization of the integral are non-negative, see ref. Sturmfels:2025wpg. (See also Appendix A.)
It is well known that Feynman integrals satisfy linear differential equations, called Picard-Fuchs equations, in their kinematic invariants, see e.g. ref. Muller-Stach:2012tgj. Alternatively, they can be viewed as first-order matrix systems of equations, as common in the high-energy physics literature on Feynman integrals, see e.g. ref. Henn:2014qga. Let us adopt the latter point of view. In the case of a single kinematic variable $x$, these equations take the following form,
@@
where ${\bf f}(x)=\{f_{1}(x),\dots,f_{k}(x)\}$ is a vector of basis integrals and $A(x)$ is a matrix. For certain natural basis choices, such as the ones we will make in this paper, the entries of $A$ are rational functions of $x$. (See e.g. ref. Henn:2014qga for more details.) In the following, we will choose basis integrals ${\bf f}(x)$ that are CM. This is expected to be always possible, as we will illustrate in section 2.4. In fact, as we explain in that section, we can find more CM integrals than strictly needed for our present purposes.
@@ 2.2 A linear program for bootstrapping Feynman integrals
It is worth emphasizing that for our purpose, it is not required to find particularly simple, or e.g. canonical Henn:2013pwa, representations of the differential equations matrix. This means that there are few prerequisites, other than availability of some form of the differential equations matrix. Also, note that if scalar CM integrals are chosen as a basis, this means that the entries of the differential equations matrix are rational, which is advantageous in view of computer algebra.
@@ 2.3.1 Differential equations
The integral $f(x)$ can be evaluated in the Feynman parametrization, cf. chapter 4 of ref. Badger:2023eqz. The result is
@@ 2.4 Multi-loop applications
In this section, we discuss applications of the method to the equal-mass ‘banana’ integrals, see Fig. 5(a), up to four loops. These integrals are relevant to phenomenology as subtopologies for processes with internal massive propagators. They provide an optimal playground for a numerical bootstrap, as they are analytically complicated with an easy to study differential equation and convergence behaviour. They are known to be associated with elliptic curves and other Calabi-Yau manifolds Bloch:2014qca, which makes their function space and analytic computation more involved. The CM bootstrap does not depend on the analytic complexity of the function space, and efficiently provides numerical values also for these more complex functions. The differential equation for these integrals is known at all loop orders and was solved up to six loops, see refs. Lairez:2022zkj; Pogel:2022vat. We demonstrate the efficient numerical evaluation of the equal-mass ‘banana’ integral families up to four loops. We discuss the runtime for different precision goals and compare the performance to other commonly used methods for numerical evaluation.
@@ 2.4.1 Choice of a suitable CM basis
We work in $D=2$ spacetime dimensions and use higher powers of propagators for sectors with several basis integrals. Finite scalar Feynman integrals are CM in the Euclidean region, as shown in ref. Henn:2024qwe. Therefore we choose basis integrals that are ultraviolet (UV) and infrared (IR) finite: The UV behaviour of a scalar $L$-loop integral in $D$ spacetime dimensions with propagator powers $\{\nu_{i}\}$ is governed by the superficial degree of divergence
@@ 2.4.2 CM bootstrap and results
Having determined a set of suitable basis integrals, we can now perform the CM bootstrap as introduced in Section 2.3.2. For each of the families, we first compute the differential equation as explained in Section 2.3.1. We use FiniteFlow Peraro:2019svx in combination with Kira Maierhofer:2017gsa to obtain and solve the relevant IBP relations. The resulting differential equations for each family are provided in ancillary files.
@@ 2.4.3 Performance analysis
We also compare the performance of our method against a commonly used program for the numerical evaluation of Feynman integrals. For convenience, we chose AMFlow Liu:2022chg. For the comparison, we compute the integrals at ten digit precision. This is a reasonable precision goal for various practical applications of numerical Feynman-integral evaluation, where phenomenological uncertainties (parton distribution functions, parameters, scale variation) typically dominate over the numerical integration error.
@@ 3.1 Stieltjes functions
Stieltjes functions are characterized by the following four fundamental properties (cf. Section 8.6 in ref. Bender):
@@
Conversely, these conditions are also sufficient to guarantee that a function $f(z)$ is of Stieltjes type, as discussed in ref. Bender. Stieltjes functions possess several useful analytical features. Most notably, they can be well approximated by Padé approximants. We note that the examples we consider in this work have a=0,b=1/Ra=0,~b=1/R, with $R\geq 0$, corresponding to functions that are analytic in the cut plane $\mathbb{C}\backslash[-R,-\infty)$. We assume these properties from here onward.
@@ 3.2 Padé approximation and convergence theorems
Padé approximations are often employed because they have many desirable features, cf. refs. BGM; Basdevant1972Pade. They usually have better convergence properties compared to Taylor approximations, and they allow one to analytically continue functions to obtain values outside the radius of convergence of the original series, (cf. refs. Davies:2019nhm; Czakon:2020vql and references therein for sample uses of Padé approximations in the context of Feynman integrals.).
@@
Figure 4: The domain $\mathcal{D}^{+}(\Delta)$ where Padé approximations are well-behaved. The convergence properties are discussed in detail in the main text. $R_{\text{max}}>0$ can take any value. $\Delta>0$ quantifies the distance from the cut on the negative real axis. Figure adapted from ref. BGM.
@@
    Convergence on the real axis (cf. Section 8.6 of ref. Bender.) For real arguments $x\in\mathbb{R}\backslash[-R,-\infty)$ and $x\geq x_{0}$, the following statements hold:
@@
    Poles and residues (cf. Theorem 5.2.1 in ref. BGM.) The poles of $P^{N+J}_{N}(x;x_{0})$ for $J\geq-1$ are all simple, lie on $(-\infty,-R)$ and have positive residues. In particular, one can write
@@
    Convergence in the cut plane (cf. Exercises 8.55, 8.59, 8.60 in ref. Bender.) Padé approximations have very good convergence away from the cut. In the cut plane $\mathbb{C}\backslash[-R,-\infty)$, both sequences $\{P^{N}_{N}(z;x_{0})\}$ and $\{P^{N-1}_{N}(z;x_{0})\}$ converge to the Stieltjes function $f(z)$, which has the expansion given in eq. (14), provided that the coefficients do not grow too quickly, more precisely, if $|a_{n}|=O((2n)!~C^{n})$ for a constant $C$ and for all $n\in\mathbb{N}$.
@@
    Error bounds (cf. Theorem 5.4.4 in ref. BGM and Theorem 16.2 in ref. Baker1975Essentials.) Padé approximants $P^{M+J}_{M}(z;x_{0})$ are good representations of the function $f(z)$ at points that are not too close to the branch cut. To make this precise, consider $\Delta>0$ and define the region
@@ 3.3 Proof of Stieltjes property for a class of Feynman integrals
    Comparing this to eq. (12), we see that $(Ax+B)^{-e}$ is a Stieltjes function for $0<e<1$, with $A>0,B>0$. Since the space of Stieltjes functions is convex Bender, the subsequent integrations over the Feynman parameters in eq. (20) preserve this property. Hence, $I(x)$ is a Stieltjes function.
@@
Scalar Feynman integrals, as defined in eq. (37), are Stieltjes functions, provided that $0<\sum_{i}\nu_{i}-LD/2\leq 1$, where $L$ is the loop order, $D$ is the space-time dimension, and $\nu_{i}$ are the propagator powers. This generalizes the CM property proven in ref. Henn:2024qwe. See Fig. 5 for examples of Feynman integrals that saturate the upper bound of the inequality.
@@ 3.4 Example relevant to one-loop five-particle scattering
As an example of the relevance of Stieltjes functions and their usefulness, we discuss functions arising in one-loop five-particle scattering processes, as studied in ref. Bern:1993mq. In that work, the amplitudes were expressed in terms of logarithms and dilogarithms. The authors of ref. Bern:1993mq identified certain special combinations of these functions that have nice properties, e.g. with respect to their singularity structure. We demonstrate that these combinations are closely related to Stieltjes functions. Let us consider the following functions,
@@
One can show that these are all Stieltjes functions, and hence they enjoy the properties discussed in section 3.1. It can then be readily verified that all transcendental functions appearing in ref. Bern:1993mq, in particular those appearing in eq. (5) of that work, can be expressed in terms of simple $\mathbb{Q}$-linear combinations of $f^{(i)}$ and their derivatives. Let us now focus on $f^{(1)}(z)$ defined in eq. (24), but for a complex argument $z$ in the cut plane, to illustrate the convergence properties of Padé approximations, discussed in section 3.2. To see that $f^{(1)}(z)$ is a Stieltjes function, consider its (dispersive) representation,

### v2

@@ 1 Introduction
Feynman integrals are crucial components in the computation of observables in perturbative quantum field theory, in particular for scattering amplitudes relevant in collider physics and gravitational wave physics Travaglini and others (2022); Huss et al. (2025); Buonanno et al. (2022). In recent years, the development of new mathematical structures has significantly advanced our ability to compute and understand Feynman integrals, see for instance ref. Abreu et al. (2022). However, the full analytic evaluation of multi-loop Feynman integrals remains a challenge, especially when multiple scales are involved. At the same time, there is considerable interest in numerically evaluating observables. It is therefore interesting to ask: How can we leverage analytic insights for numerical evaluation of Feynman integrals?
In the literature, a number of promising methods exist, as we review presently. These include methods based on contour deformation and sector decomposition Binoth and Heinrich (2000) of Feynman integrals in Feynman parametrization. Sector decomposition helps to disentangle overlapping singularities by reorganizing the integral into a form in which divergences become more transparent and easier to handle numerically. See for refs. Heinrich et al. (2024); Smirnov (2016) for computer implementations of this method. The authors of ref. Borinsky et al. (2023) use tropical Monte-Carlo-based sampling methods. This method, while powerful and performant, is currently available for finite integrals only.
Another approach is to derive and to numerically evaluate differential equations satisfied by Feynman integrals. For example, refs. Hidding (2021); Armadillo et al. (2023); Prisco et al. (2025); Mezzarobba (2025) rely on knowing the value of the integrals at some reference point (e.g. as determined by physical consistency conditions, or via other means, such as numerical evaluation via the above methods), and then analytically continue that value to other kinematic values by matching the domains of convergence of series expansions about different kinematic points. For this to work well, it is desirable to have the differential equation in a form as simple as possible (cf. the review Henn (2015)), which requires additional steps. On the other hand, ref. Liu and Ma (2023) introduces an auxiliary parameter, with the help of which it becomes easy to determine boundary values of an integral, which are then analytically continued to the integral without the auxiliary parameter. While very powerful and universally applicable, introducing the additional scale of the auxiliary parameter sometimes constitutes a bottleneck.
@@
The first approach we consider also takes differential equations as input, but does not require any specific form of them, which makes our approach easy to use and without complicated prerequisites. More specifically, we propose a new method for bootstrapping Feynman integrals, by combining their differential equations with the constraint of complete monotonicity. Complete monotonicity is a powerful positivity condition on (all) derivatives of a function. In ref. Henn and Raman (2025) two of the present authors showed that many relevant quantities in quantum field theory are completely monotone. See refs. Ma and Weigert (2025); Mazzucchelli and Raman (2025) for further recent research in this direction.
In particular, scalar Feynman integrals exhibit complete monotonicity in their Euclidean region with respect to kinematic invariants Henn and Raman (2025). When combined with differential equations for a set of integrals, this property imposes powerful constraints on the space of possible solutions. We show that this interplay between differential equations and complete monotonicity can uniquely determine the integral. This extends previous work on numerically bootstrapping Feynman integrals Zeng (2023), which employed positivity properties of functions, but did not consider derivatives.
@@
Stieltjes functions enjoy remarkable analyticity properties that imply that they can be extremely well described by Padé approximants Basdevant (1972); Baker and Graves-Morris (1996). In particular, strong bounds and convergence theorems are known for the Padé approximants. Leveraging this knowledge, we show how one can obtain reliable numerical approximations over a wide range of kinematic space, including complex-valued, analytically continued results in physical regions. This allows for a fast and high-precision numerical determination of integrals across a wide range of kinematics by exploiting minimal analytic assumptions to obtain tight numerical constraints.
@@ 2.1 Complete monotonicity and Feynman integrals
A real-valued function $f:R\to\mathbb{R}$ is said to be completely monotonic (CM) in the region $R$ if it satisfies (cf. ref. Widder (1941))
@@
There is a straightforward connection to Feynman integrals that was uncovered in ref. Henn and Raman (2025). In that reference, two of the present authors showed that scalar Feynman integrals are CM functions of appropriate linear combinations of kinematic variables, within their Euclidean region. In this context, we understand by Euclidean region the set of kinematic values for which the graph polynomials appearing in Feynman parametrization of the integral are non-negative, see ref. Sturmfels and Telek (2025). (See also Appendix A.)
It is well known that Feynman integrals satisfy linear differential equations, called Picard-Fuchs equations, in their kinematic invariants, see e.g. ref. Müller-Stach et al. (2014). Alternatively, they can be viewed as first-order matrix systems of equations, as common in the high-energy physics literature on Feynman integrals, see e.g. ref. Henn (2015). Let us adopt the latter point of view. In the case of a single kinematic variable $x$, these equations take the following form,
@@
where ${\bf f}(x)=\{f_{1}(x),\dots,f_{k}(x)\}$ is a vector of basis integrals and $A(x)$ is a matrix. For certain natural basis choices, such as the ones we will make in this paper, the entries of $A$ are rational functions of $x$. (See e.g. ref. Henn (2015) for more details.) In the following, we will choose basis integrals ${\bf f}(x)$ that are CM. This is expected to be always possible, as we will illustrate in section 2.4. In fact, as we explain in that section, we can find more CM integrals than strictly needed for our present purposes.
@@ 2.2 A linear program for bootstrapping Feynman integrals
It is worth emphasizing that for our purpose, it is not required to find particularly simple, or e.g. canonical Henn (2013), representations of the differential equations matrix. This means that there are few prerequisites, other than availability of some form of the differential equations matrix. Also, note that if scalar CM integrals are chosen as a basis, this means that the entries of the differential equations matrix are rational, which is advantageous in view of computer algebra.
@@ 2.3.1 Differential equations
The integral $f(x)$ can be evaluated in the Feynman parametrization, cf. chapter 4 of ref. Badger et al. (2024). The result is
@@ 2.4 Multi-loop applications
In this section, we discuss applications of the method to the equal-mass ‘banana’ integrals, see Fig. 5(a), up to four loops. These integrals are relevant to phenomenology as subtopologies for processes with internal massive propagators. They provide an optimal playground for a numerical bootstrap, as they are analytically complicated with an easy to study differential equation and convergence behaviour. They are known to be associated with elliptic curves and other Calabi-Yau manifolds Bloch et al. (2015), which makes their function space and analytic computation more involved. The CM bootstrap does not depend on the analytic complexity of the function space, and efficiently provides numerical values also for these more complex functions. The differential equation for these integrals is known at all loop orders and was solved up to six loops, see refs. Lairez and Vanhove (2023); Pögel et al. (2023). We demonstrate the efficient numerical evaluation of the equal-mass ‘banana’ integral families up to four loops. We discuss the runtime for different precision goals and compare the performance to other commonly used methods for numerical evaluation.
@@ 2.4.1 Choice of a suitable CM basis
We work in $D=2$ spacetime dimensions and use higher powers of propagators for sectors with several basis integrals. Finite scalar Feynman integrals are CM in the Euclidean region, as shown in ref. Henn and Raman (2025). Therefore we choose basis integrals that are ultraviolet (UV) and infrared (IR) finite: The UV behaviour of a scalar $L$-loop integral in $D$ spacetime dimensions with propagator powers $\{\nu_{i}\}$ is governed by the superficial degree of divergence
@@ 2.4.2 CM bootstrap and results
Having determined a set of suitable basis integrals, we can now perform the CM bootstrap as introduced in Section 2.3.2. For each of the families, we first compute the differential equation as explained in Section 2.3.1. We use FiniteFlow Peraro (2019) in combination with Kira Maierhöfer et al. (2018) to obtain and solve the relevant IBP relations. The resulting differential equations for each family are provided in ancillary files.
@@ 2.4.3 Performance analysis
We also compare the performance of our method against a commonly used program for the numerical evaluation of Feynman integrals. For convenience, we chose AMFlow Liu and Ma (2023). For the comparison, we compute the integrals at ten digit precision. This is a reasonable precision goal for various practical applications of numerical Feynman-integral evaluation, where phenomenological uncertainties (parton distribution functions, parameters, scale variation) typically dominate over the numerical integration error.
@@ 3.1 Stieltjes functions
Stieltjes functions are characterized by the following four fundamental properties (cf. Section 8.6 in ref. Bender and Orszag (1999)):
@@
Conversely, these conditions are also sufficient to guarantee that a function $f(z)$ is of Stieltjes type, as discussed in ref. Bender and Orszag (1999). Stieltjes functions possess several useful analytical features. Most notably, they can be well approximated by Padé approximants. We note that the examples we consider in this work have a=0,b=1/Ra=0,~b=1/R, with $R\geq 0$, corresponding to functions that are analytic in the cut plane $\mathbb{C}\backslash[-R,-\infty)$. We assume these properties from here onward.
@@ 3.2 Padé approximation and convergence theorems
Padé approximations are often employed because they have many desirable features, cf. refs. Baker and Graves-Morris (1996); Basdevant (1972). They usually have better convergence properties compared to Taylor approximations, and they allow one to analytically continue functions to obtain values outside the radius of convergence of the original series, (cf. refs. Davies et al. (2019); Czakon and Niggetiedt (2020) and references therein for sample uses of Padé approximations in the context of Feynman integrals.).
@@
Figure 4: The domain $\mathcal{D}^{+}(\Delta)$ where Padé approximations are well-behaved. The convergence properties are discussed in detail in the main text. $R_{\text{max}}>0$ can take any value. $\Delta>0$ quantifies the distance from the cut on the negative real axis. Figure adapted from ref. Baker and Graves-Morris (1996).
@@
    Convergence on the real axis (cf. Section 8.6 of ref. Bender and Orszag (1999).) For real arguments $x\in\mathbb{R}\backslash[-R,-\infty)$ and $x\geq x_{0}$, the following statements hold:
@@
    Poles and residues (cf. Theorem 5.2.1 in ref. Baker and Graves-Morris (1996).) The poles of $P^{N+J}_{N}(x;x_{0})$ for $J\geq-1$ are all simple, lie on $(-\infty,-R)$ and have positive residues. In particular, one can write
@@
    Convergence in the cut plane (cf. Exercises 8.55, 8.59, 8.60 in ref. Bender and Orszag (1999).) Padé approximations have very good convergence away from the cut. In the cut plane $\mathbb{C}\backslash[-R,-\infty)$, both sequences $\{P^{N}_{N}(z;x_{0})\}$ and $\{P^{N-1}_{N}(z;x_{0})\}$ converge to the Stieltjes function $f(z)$, which has the expansion given in eq. (14), provided that the coefficients do not grow too quickly, more precisely, if $|a_{n}|=O((2n)!~C^{n})$ for a constant $C$ and for all $n\in\mathbb{N}$.
@@
    Error bounds (cf. Theorem 5.4.4 in ref. Baker and Graves-Morris (1996) and Theorem 16.2 in ref. Baker Jr. (1975).) Padé approximants $P^{M+J}_{M}(z;x_{0})$ are good representations of the function $f(z)$ at points that are not too close to the branch cut. To make this precise, consider $\Delta>0$ and define the region
@@ 3.3 Proof of Stieltjes property for a class of Feynman integrals
    Comparing this to eq. (12), we see that $(Ax+B)^{-e}$ is a Stieltjes function for $0<e<1$, with $A>0,B>0$. Since the space of Stieltjes functions is convex Bender and Orszag (1999), the subsequent integrations over the Feynman parameters in eq. (20) preserve this property. Hence, $I(x)$ is a Stieltjes function.
@@
Scalar Feynman integrals, as defined in eq. (37), are Stieltjes functions, provided that $0<\sum_{i}\nu_{i}-LD/2\leq 1$, where $L$ is the loop order, $D$ is the space-time dimension, and $\nu_{i}$ are the propagator powers. This generalizes the CM property proven in ref. Henn and Raman (2025). See Fig. 5 for examples of Feynman integrals that saturate the upper bound of the inequality.
@@ 3.4 Example relevant to one-loop five-particle scattering
As an example of the relevance of Stieltjes functions and their usefulness, we discuss functions arising in one-loop five-particle scattering processes, as studied in ref. Bern et al. (1993). In that work, the amplitudes were expressed in terms of logarithms and dilogarithms. The authors of ref. Bern et al. (1993) identified certain special combinations of these functions that have nice properties, e.g. with respect to their singularity structure. We demonstrate that these combinations are closely related to Stieltjes functions. Let us consider the following functions,
@@
One can show that these are all Stieltjes functions, and hence they enjoy the properties discussed in section 3.1. It can then be readily verified that all transcendental functions appearing in ref. Bern et al. (1993), in particular those appearing in eq. (5) of that work, can be expressed in terms of simple $\mathbb{Q}$-linear combinations of $f^{(i)}$ and their derivatives. Let us now focus on $f^{(1)}(z)$ defined in eq. (24), but for a complex argument $z$ in the cut plane, to illustrate the convergence properties of Padé approximations, discussed in section 3.2. To see that $f^{(1)}(z)$ is a Stieltjes function, consider its (dispersive) representation,